OpenStudy (_.mufasa._):

Solve for x. State any restrictions on the variables. a(bx+2)=cx-12; A. x=(-12-2a)/(ab-c);ab≠c B. x=(-12+2a)/(c+ab);ab≠c C. x=(2a+12)/(c-ab);no restrictions D. x=(ab-c)/(-12-2a);a≠-1

OpenStudy (campbell_st):

well distribute on the left side abx + 2a = cx - 12 collect like terms abx - cx = -2a - 12 factor x(ab-c) = (-2a -12) divide both sides by (ab - c) that will be the solution

OpenStudy (campbell_st):

then when you get a solution... have a look at the denominator... and decide if there are restrictions

OpenStudy (anonymous):

x=- 2(a+6)/ab-c

OpenStudy (_.mufasa._):

@campbell_st so its A. x -12 -2a\[x=\frac{ -12-2a }{ ab-c } ; ab \neq c \] ?

OpenStudy (campbell_st):

that's correct, the order in the numerator is reversed to what I wrote when I solved it

OpenStudy (_.mufasa._):

@campbell_st thank you!